Find π‘Μ‚π‘Žπ‘›π‘‘ π‘žΜ‚, given X and n.







1. X = 56, n = 80

1
Expert's answer
2022-05-30T22:57:41-0400
"\hat{p}=\dfrac{X}{n}=\dfrac{56}{80}=0.7"

"\hat{q}=1-\hat{p}=1-0.7=0.3"